OpenStudy (anonymous):

Find The dISTANCE Between Each Pair Of Points: (0,2 & 1/6) , (3, & 1/3)\

OpenStudy (anonymous):

1. sqrt(x2-x1)^2+(y2-y1)^2

OpenStudy (anonymous):

1.sqt(1/3-1/6)^2+(3-0)^2

OpenStudy (anonymous):

sqrt(1/6)^2+(3)^2 = sqrt(1/36+9) =3